One number is 2 times another. If the sum of their reciprocal is 1/4, find the two numbers.

Thank You!!!!

Answer by stanbon(48558)

One number is 2 times another. If the sum of their reciprocal is 1/4, find the two numbers.
Let one of the numbers be "x"
The other is "2x"
EQUATION:
1/x + 1/2x = 1/4
(2x+x)/2x^2 = 1/4
3x/2x^2 = 1/4
3/2x=1/4
2x=12 (one of the numbers)
x=6 (the other number)
Cheers,
Stan H.